# Self-Exploring Language Models: Active Preference Elicitation for Online Alignment

### v1

@@ Appendix C Experiment Setup
In experiments, we use the Alignment Handbook [56] framework as our codebase. We find the best hyperparameter settings by conducting a grid search over the iteration number, batch size, learning rate, and label update rule for the iterative DPO baseline. The results for the Zephyr-based models are shown in Figure 5. Specifically, we find that using the same amount of data, updating the model too many iterations can lead to instability. So we set the iteration number to $3$ for Llama3-It-based and Zephyr-based models (excluding the first iteration of DPO training). Besides, we observe that choosing different batch sizes has a large effect on the models’ performance and the optimal batch size heavily depends on the model architecture. In experiments, we set the batch size to $256$ and $128$ for the Zephyr-based and Llama3-It-based models, respectively. For the learning rate, we consider three design choices: cyclic learning rate with constant cycle amplitude, linearly decayed cycle amplitude, and decayed cycle amplitude at the last iteration. We find that a decaying cycle amplitude performs better than constant amplitudes in general. Thus, for Zephyr-based models, we set the learning to $75e-$ 7 for the first three iterations and $71e-$ 7 for the last iteration. In each iteration, the warmup ratio is $0.1$. For Llama3-It-based models, we use a linearly decayed learning rate from $75e-$ 7 to $71e-$ 7 within $3$ iterations with the same warmup ratio. We also test two update ways for the preference data. One is to rank $y_{w},y_{l},y_{\text{ref}}$ and keep the best and worst responses in the updated dataset, which is the setting that is described in the main paper. The other is to compare $y_{w}$ and $y_{\text{ref}}$ and replace the chosen or rejected response by $y_{\text{ref}}$ based on the comparison result. We find that the former design performs better than the latter. We also compared with stepwise DPO [28], which updates the reference model at each iteration but uses the original dataset instead of the updated one. This demonstrates that exploring and collecting new data is necessary.

### v3

@@ Appendix D Experiment Setup
In experiments, we use the Alignment Handbook (Tunstall et al., 2023a) framework as our codebase. We find the best hyperparameter settings for the strong iterative DPO baseline by conducting a grid search over the iteration number, batch size, learning rate, and label update rule. The results for the Zephyr-based models are shown in Figure 6. Specifically, we find that using the same amount of data, updating the model too many iterations can lead to instability. So we set the iteration number to $3$ for Llama3-It-based and Zephyr-based models (excluding the first iteration of DPO training). Besides, we observe that choosing different batch sizes has a large effect on the models’ performance and the optimal batch size heavily depends on the model architecture. In experiments, we set the batch size to $256$ and $128$ for the Zephyr-based and Llama3-It-based models, respectively. For the learning rate, we consider three design choices: cyclic learning rate with constant cycle amplitude, linearly decayed cycle amplitude, and decayed cycle amplitude at the last iteration. We find that a decaying cycle amplitude performs better than constant amplitudes in general. Thus, for Zephyr-based models, we set the learning to $75e-$ 7 for the first three iterations and $71e-$ 7 for the last iteration. In each iteration, the warmup ratio is $0.1$. For Llama3-It-based models, we use a linearly decayed learning rate from $75e-$ 7 to $71e-$ 7 within $3$ iterations with the same warmup ratio. We also test two update ways for the preference data. One is to rank $y_{w},y_{l},y_{\text{ref}}$ and keep the best and worst responses in the updated dataset, which is the setting that is described in the main paper. The other is to compare $y_{w}$ and $y_{\text{ref}}$ and replace the chosen or rejected response by $y_{\text{ref}}$ based on the comparison result. We find that the former design performs better than the latter. We also compared with stepwise DPO (Kim et al., 2024), which updates the reference model at each iteration but uses the original dataset instead of the updated one. This demonstrates that exploring and collecting new data is necessary.